OpenStudy (anonymous):

Simplify 16 / 2 + 6 – 4 x 3

OpenStudy (anonymous):

16 ÷ 2 = 8 8 + 6 = 14 14 - 4 = 10 10 • 3 =? Just do it one step at a time :)

OpenStudy (lgbasallote):

careful there @rebeccaskell94 ...try to remember you PEMDAS ;)

OpenStudy (lgbasallote):

your*

OpenStudy (anonymous):

30

OpenStudy (lgbasallote):

it's not 30....try to redo this (16/2) + 6 - (4x3)

OpenStudy (lgbasallote):

remember to multiply 4 by 3 before subtracting according to PEMDAS

OpenStudy (anonymous):

12

OpenStudy (lgbasallote):

that is 4 x 3

OpenStudy (lgbasallote):

what's the rest?

OpenStudy (lgbasallote):

16/2 = 8 8 + 6 = 14 14 - (4x3) = 14 - 12

OpenStudy (anonymous):

16 divide by 2? which is 8

OpenStudy (lgbasallote):

yes 16 divided by 2 is 8

OpenStudy (lgbasallote):

then you add 6 it becomes 14

OpenStudy (anonymous):

20

OpenStudy (lgbasallote):

then yousubtract (4x3) 4 x 3 is 12 so you subtract 12 from 14 what is 14-12

OpenStudy (lgbasallote):

will you please stop enlisting the multiple choice and actually listen :p

OpenStudy (lgbasallote):

what is 14 - 12

OpenStudy (anonymous):

2

OpenStudy (lgbasallote):

then that's the answer

OpenStudy (anonymous):

I am listening...........

OpenStudy (lgbasallote):

well it's already over lol

OpenStudy (lgbasallote):

14 - 12 = 2 therefore that's the answer